# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v3

@@ Abstract
Recent advances in neural surrogate modeling offer the potential for transformative innovations in applications such as automotive aerodynamics. Yet, industrial-scale problems often involve volumetric meshes with cell counts reaching 100 million, presenting major scalability challenges. Complex geometries further complicate modeling through intricate surface-volume interactions, while quantities such as vorticity are highly nonlinear and must satisfy strict divergence-free constraints. To address these requirements, we introduce  Anchored-Branched Universal Physics Transformers (AB-UPT) as a novel modeling scheme for building neural surrogates for  computational fluid dynamics (CFD) simulations. AB-UPT is designed to: (i) decouple geometry encoding and prediction tasks via multi-branch operators; (ii) enable scalability to high-resolution outputs via neural simulation in a low-dimensional latent space, coupled with anchored neural field decoders to predict high-fidelity outputs; (iii) enforce physics consistency by a novel divergence-free formulation. We show that AB-UPT yields state-of-the-art predictive accuracy of surface and volume fields on automotive CFD simulations ranging from 33 thousand up to 150 million mesh cells. Furthermore, our anchored neural field architecture enables the enforcement of hard physical constraints on the physics predictions without degradation in performance, exemplified by modeling divergence-free vorticity fields. Notably, the proposed models can be trained on a single GPU in less than a day and predict industry-standard surface and volume fields within seconds. Additionally, we show that the flexible design of our method enables neural simulation from a computer-aided design geometry alone, omitting the need for costly CFD meshing procedures.
@@ 1 Introduction
Figure 1: AB-UPT is a neural surrogate model trained to jointly model surface and volume variables of automotive CFD simulations with $>100$M simulation mesh cells. It obtains state-of-the-art surface and volume predictions (left), accurately models drag and lift coefficients (center), all on a single GPU (right).
@@
Scaling large machine learning models is usually framed in terms of the number of trainable parameters and the overall dataset size. Particular to the CFD field, however, is the fact that industry-standard volumetric meshes often exceed more than 100 million cells (i. e., high-dimensional data), while at the same time the amount of simulation samples is low (i. e., low number of independent training samples). Moreover, intricate geometry-dependent interactions between the surface and volumetric fields govern the underlying physics. Further complexity arises from field quantities such as vorticity, which are highly nonlinear yet inherently divergence-free. The combination of high-dimensional non-linear physics phenomena and the low number of industry-standard training samples makes scaling neural surrogate models for CFD particularly hard. Thus, our definition of scale also encompasses problem scale, i. e. the fidelity of the simulation (visually represented in Figure 2).
@@
To harness problem scale, we propose Anchored-Branched Universal Physics Transformers (AB-UPT) (see Figure 3, a novel modeling scheme, build on the Universal Physics Transformer (UPT) framework, for building neural surrogates for automotive CFD simulations. The modeling is organized into three branches (geometry-encoding, surface-prediction, volume-prediction), with cross-attention between and self-attention within branches, decoupling feature extraction from field prediction. Within each branch, the encoded simulation points are partitioned into two sets: a small set of selected anchor tokens share information among each other via self-attention and together provide context (e. g., about the physical dynamics); the remaining query tokens only receive information from the anchor tokens via cross-attention, leading to independent query predictions. This anchor-based design has three critical advantages, resulting in strong generalization performance, inference efficiency, and physics consistency: (i) selecting subsets of points massively increases the amount of effective training instances, and the number of anchor tokens governs the model’s computational capacity; (ii) computational cost scales quadratically only with the small number of anchor tokens, while predictions on all other points have linear computational complexity; (iii) interpreting the anchor mechanism as a conditional-field model enables the application of differential operators, and we leverage this capability to derive a divergence-free vorticity-field formulation.
@@ 2 Preliminaries and related work
HRLES Equations Numeric - - ✗✗ -
RANS Equations Numeric * - - -
@@
AB-UPT Transformer Latent space
Table 1: Conceptual overview of different CFD solvers. Numeric integration schemes, while accurate, are slow and require specialized meshes to converge, with several orders of magnitude variations in fidelity and runtime. Neural surrogates struggle to be trained on the full-resolution mesh, which makes them slow to train or requires multi-GPU setups for million-scale meshes. Next to that, those methods require computation units that favor speed over accuracy. We introduce AB-UPT, a new method incorporating anchor tokens. AB-UPT unlocks linear problem scaling, which enables training of neural surrogates on a single GPU in less than a day, even on million-scale meshes. This efficiency also allows us to leverage expressive neural computation units such as the Transformer. Furthermore, our multi-branch design enables simulation with any input mesh, uses reduced latent modeling of physical dynamics. Finally, anchor attention acts as a neural field, enabling an efficient divergence-free vorticity field formulation. *It’s difficult to make comparisons between RANS simulations and models that are trained on HRLES simulations, since RANS simulations cannot resolve certain phenomena which HRLES can resolve. One example is strong separation.
@@ 2.1 CFD for automotive aerodynamics
are used (Ashton et al., 2024b), where $\bm{e}_{flow}$ is a unit vector into the free stream direction, $\bm{e}_{lift}$ a unit vector into the lift direction perpendicular to the free stream direction, $\rho$ the density, $v$ the free stream velocity, and $A_{ref}$ a characteristic reference area. Predicting these surface integrals allows for an efficient estimation when using deep learning surrogates, since these surrogates can directly predict the surface values without the need to model the full 3D volume field, as required by numerical CFD simulations.
@@ 2.2 Transformer blocks for building neural surrogates
Neural fields (Mildenhall et al., 2020) are neural networks that take continuous coordinates as inputs and produce a prediction for this exact coordinate. The continuous nature of the input coordinates enables evaluating this function at arbitrary positions, making it a field. By using embedded coordinates as $\mathbf{Z}$ and a latent representation of, e.g., the simulation state as $\mathbf{Z}_{KV}$, cross-attention acts as a conditional neural field (Wang et al., 2024b) as each query vector retrieves information from the latent representation to produce a prediction for a given query coordinate. It is worth noting that the neural field decoding via cross-attention operates on a point-wise basis, and hence, is independent of the number of query points.
@@ 3 Anchored-Branched Universal Physics Transformers
We introduce Anchored-Branched Universal Physics Transformers (AB-UPT), which leverages two key modeling components: (i) the multi-branch architecture outlined in LABEL:{sec:method-multibranch} separates geometry encoding, surface-level simulation, and volume-level simulation into distinct branches that interact with each other; (ii) anchor attentiondescribed in Section 3.2 enables modeling of complex physical dynamics by leveraging expressive dot-product attention (Vaswani et al., 2017) within a heavily reduced set of anchor tokens. To generate predictions for the full simulation mesh during inference, anchor attention employs cross-attention between (millions) of query tokens and the small amount of anchor tokens, which has linear inference complexity. Additionally, anchor attention can be seen as a conditional neural field, which allows creating arbitrary output resolutions, and enables the application of differential operators to enforce hard physical constraints such as divergence-free vorticity-field predictions (see Section 3.3).
@@ 3.1 Multi-branch architecture
Figure 4: The Anchored-Branched Universal Physics Transformers model is a multi-branch Transformer neural operator using three distinct branches: a surface, geometry, and volume branch. The physics blocks within each branch share parameters and interact through cross-attention. To handle problems at industry-scale, we select a subset of $M$ points from the input point cloud. We call these points anchor tokens, and they serve as the keys and values for attention computation. This approach significantly reduces the complexity of attention, making it possible to scale our model to industry-standard simulation meshes.
Our model builds on the Universal Physics Transformer (UPT(Alkin et al., 2024a) and multi-branch Transformers for neural operators (Alkin et al., 2024b; Esser et al., 2024). Multi-branch Transformers consist of multiple branches (which can share parameters) that each handle different data streams or modalities. We define a geometry branch, a surface branch, and a volume branch (see Figures 3 and 4 for an overview). Those branches interact differently in respective encoding, physics, and decoding blocks, and allow us to enable scalability to high-resolution outputs via the new anchor attention. We distinguish the geometry, surface, and volume branches by the superscripts ${{g}}$, ${{s}}$, and ${{v}}$, respectively.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The surface and volume branch consists of $K$ physics blocks. Each physics block $k$, where $k\in\{1,\dots,K\}$, consists of a self-attention and cross-attention block, which share parameters between the surface and volume branches. Each physics block starts with a self-attention block, where the queries, keys, and values are all derived from the output $\mathbf{Z}_{k-1}$ of the preceding physics block $k-1$, where $\mathbf{Z}_{0}$ from Equation 7 is the input to the first block. Next is a cross-attention block, enabling surface-volume interaction. For the surface branch, the queries originate from the previous block within the surface branch. However, the keys and values come from the volume branch $k$. For the volume branch, however, it is reversed: the queries come from the volume branch, and the keys and values from the surface branch. These cross-attention layers allow for intricate interactions between the surface and volumetric fields. Formally, physics block $k$ for the surface and volume branch, $\mathcal{B}^{{s}}_{k}$ and $\mathcal{B}^{{v}}_{k}$, are defined by:
[…]
@@ 3.2 Anchor attention
High-fidelity numerical simulations in industry often rely on extremely fine mesh discretizations, sometimes comprising hundreds of millions of cells (Ashton et al., 2024b). In practice, representing such large point clouds (meshes) as inputs to surrogate models becomes computationally challenging. In particular, Transformer models (Vaswani et al., 2017)—the workhorse of many modern deep learning architectures—capture dependencies among all $N$ input tokens via self-attention:
@@
To mitigate this problem, we introduce the concept of anchor attention, which splits the discretized input mesh (i. e., the point cloud) into two parts: (i) a small subset $\mathbf{A}$ of $M$ anchor tokens is uniformly sampled from the $N$ input tokens and used to derive the key matrix $\mathbf{K}(\mathbf{A})$ and value matrix $\mathbf{V}(\mathbf{A})$ for the attention mechanism; (ii) the query matrix $\mathbf{Q}(\mathbf{Z})$, which is derived from all $N$ input tokens, then cross-attend to the anchor key and value matrices. Intuitively, this formulation splits the $N$ input tokens into $M$ anchor tokens (used as queries, keys, and values in the attention) and $N-M$ query tokens (used only as queries in the attention). It is equivalent to self-attention within the $M$ anchor tokens and cross-attention between $N-M$ query tokens and the $M$ anchor tokens. Formally, anchor attention is defined as follows:
@@
This design allows us to focus on problem scale (c. f. Figure 2): by choosing the number of anchor tokens $M$, we directly control the model’s computational capacity, while enabling pointwise predictions at arbitrary points in the continuous output domain. These anchor tokens span both the surface and volume mesh, and their selection is crucial for capturing problem-specific structure. Increasing $M$ enhances contextual information available at each query location without changing the amount of model parameters, thereby improving performance up to a saturation point (see Figure 5).
@@
Importantly, the $N-M$ query tokens only need to compute attention with a much smaller set of $M$ key-value pairs (derived from the anchor tokens) compared to full self-attention, which computes all $N^{2}$ pairs. Anchor attention therefore reduces the computational complexity of the attention layers from $\mathcal{O}\left(N^{2}\right)$ to $\mathcal{O}\left(M^{2}+(N-M)M\right)=\mathcal{O}\left(MN\right)$, where $M^{2}$ is the self-attention between the $M$ anchor tokens and $(N-M)M$ is the cross-attention from $N-M$ query tokens to $M$ anchor tokens. This is crucial because surrogate models often need to be evaluated at a large number of input points (e. g., to obtain accurate drag/lift coefficient predictions as shown in Figure 1).
@@
Additionally, since there are no interactions between query tokens, an arbitrary number of query tokens can be used per forward pass. Therefore, query tokens can be processed in chunks, enabling constant memory consumption by using a constant number of queries. When processing all $N$ input tokens at once (end-to-end inference), the memory complexity of anchor attention is $\mathcal{O}(N)$ (we use an efficient attention implementation (Dao et al., 2022)). However, as $N$ can be $>$ 100 million, it would still run into out-of-memory errors on a single GPU. By chunking the $N-M$ query tokens into chunks of size $C$, memory complexity reduces to $\mathcal{O}(M+C)$, i.e., constant memory consumption. We refer to this methodology as chunked inference (c. f. inference memory in Figure 1). Moreover, in chunked inference, the key-value pairs of anchor tokens can be cached when calculating them in the first chunked forward pass to omit recalculating them for subsequent chunks (anchor tokens remain constant for all chunks). However, as $M\ll N$, this is often negligible.
@@
Finally, by treating the anchors as conditioning set for the queries, anchor attention turns the architecture into a conditional field model: the field values at any query location are predicted from the query location itself and the context provided by the anchors, but independently from any other query location. This has two important benefits. First, this enables creating predictions at arbitrary query positions, which allows AB-UPT to predict at arbitrary resolution. Many simulation insights can be obtained from much lower resolutions than necessary for numeric simulation (e.g., drag/lift coefficients in Figure 6). Second, it allows us to impose physics-based constraints directly into the model predictions, such as a divergence-free condition for vorticity (see Section 3.3). In the context of Neural Processes, the concept of conditioning the output prediction on a (small) subset of context has been explored in (Feng et al., 2023; Ashman et al., 2024).

### v4

@@ Abstract
Recent advances in neural surrogate modeling offer the potential for transformative innovations in applications such as automotive aerodynamics. Yet, industrial-scale problems often involve volumetric meshes with cell counts reaching 100 million, presenting major scalability challenges. Complex geometries further complicate modeling through intricate surface-volume interactions, while quantities such as vorticity are highly nonlinear and must satisfy strict divergence-free constraints. To address these requirements, we introduce  Anchored-Branched Universal Physics Transformers (AB-UPT) as a novel modeling scheme for building neural surrogates for  computational fluid dynamics (CFD) simulations. AB-UPT is designed to: (i) decouple geometry encoding and prediction tasks via multi-branch operators; (ii) enable scalability to high-resolution outputs via neural simulation in a low-dimensional latent space, coupled with anchored neural field decoders to predict high-fidelity outputs; (iii) enforce physics consistency by a divergence-free formulation. We show that AB-UPT yields state-of-the-art predictive accuracy of surface and volume fields on automotive CFD simulations ranging from 33 thousand up to 150 million mesh cells. Furthermore, our anchored neural field architecture enables the enforcement of hard physical constraints on the physics predictions without degradation in performance, exemplified by modeling divergence-free vorticity fields. Notably, the proposed models can be trained on a single GPU in less than a day and predict industry-standard surface and volume fields within seconds. Additionally, we show that the flexible design of our method enables neural simulation from a computer-aided design geometry alone, thereby eliminating the need for costly CFD meshing procedures for inference.
@@ 1 Introduction
Figure 1: AB-UPT is a neural surrogate model trained to jointly model surface and volume variables of automotive CFD simulations with >100>100M simulation mesh cells. It obtains state-of-the-art surface and volume predictions (left), accurately models drag and lift coefficients (center), all on a single GPU (right).
@@
Scaling large machine learning models is usually framed in terms of the number of trainable parameters and the overall dataset size. Particular to the CFD field, however, is the fact that industry-standard volumetric meshes often exceed more than 100 million cells (i. e., high-dimensional data), while at the same time the amount of simulation samples is low (i. e., low number of independent training samples). Moreover, intricate geometry-dependent interactions between the surface and volumetric fields govern the underlying physics. Further complexity arises from field quantities such as vorticity, which are highly nonlinear yet inherently divergence-free. The combination of high-dimensional non-linear physics phenomena and the low number of industry-standard training samples makes scaling neural surrogate models for CFD particularly hard. Thus, our definition of scale also encompasses problem scale, i. e. the fidelity of the simulation (visually represented in Figure˜2).
@@
To harness problem scale, we propose Anchored-Branched Universal Physics Transformers (AB-UPT) (see Figure˜3, a novel modeling scheme, build on the Universal Physics Transformer (UPT) framework, for building neural surrogates for automotive CFD simulations. The modeling is organized into three branches (geometry-encoding, surface-prediction, volume-prediction), with cross-attention between and self-attention within branches, decoupling feature extraction from field prediction. Within each branch, the encoded simulation points are partitioned into two sets: a small set of selected anchor tokens share information among each other via self-attention and together provide context (e. g., about the physical dynamics); the remaining query tokens only receive information from the anchor tokens via cross-attention, leading to independent query predictions. This anchor-based design has three critical advantages, resulting in strong generalization performance, inference efficiency, and physics consistency: (i) selecting subsets of points massively increases the amount of effective training instances, and the number of anchor tokens governs the model’s computational capacity; (ii) computational cost scales quadratically only with the small number of anchor tokens, while predictions on all other points have linear computational complexity; (iii) interpreting the anchor mechanism as a conditional‐field model enables the application of differential operators, and we leverage this capability to derive a divergence-free vorticity-field formulation.
@@ 2 Preliminaries and related work
HRLES Equations Numeric ✓✓✓ - - ✗✗ -
RANS Equations Numeric ✓✓* - - -
@@
AB-UPT Transformer Latent space ✓✓
Table 1: Conceptual overview of different CFD solvers. Numeric integration schemes, while accurate, are slow and require specialized meshes to converge, with several orders of magnitude variations in fidelity and runtime. Neural surrogates struggle to be trained on the full-resolution mesh, which makes them slow to train or requires multi-GPU setups for million-scale meshes. Next to that, those methods require computation units that favor speed over accuracy. We introduce AB-UPT, a new method incorporating anchor tokens. AB-UPT unlocks linear problem scaling, which enables training of neural surrogates on a single GPU in less than a day, even on million-scale meshes. This efficiency also allows us to leverage expressive neural computation units such as the Transformer. Furthermore, our multi-branch design enables simulation with any input mesh, uses reduced latent modeling of physical dynamics. Finally, anchor attention acts as a neural field, enabling an efficient divergence-free vorticity field formulation. *It’s difficult to make comparisons between RANS simulations and models that are trained on HRLES simulations, since RANS simulations cannot resolve certain phenomena that HRLES can resolve. One example is strong separation.
@@ 2.1 CFD for automotive aerodynamics
are used (Ashton et al., 2024b), where $\bm{e}_{\text{flow}}$ is a unit vector into the free stream direction, $\bm{e}_{\text{lift}}$ a unit vector into the lift direction perpendicular to the free stream direction, $\rho$ the density, $v$ the magnitude of the free stream velocity, and $A_{\text{ref}}$ a characteristic reference area. Predicting these surface integrals allows for an efficient estimation when using deep learning surrogates, since these surrogates can directly predict the surface values without the need to model the full 3D volume field, as required by numerical CFD simulations.
@@ 2.2 Transformer blocks for building neural surrogates
Neural fields (Mildenhall et al., 2020) are neural networks that take continuous coordinates as inputs and produce a prediction for this exact coordinate. The continuous nature of the input coordinates enables evaluating this function at arbitrary positions, making it a field. By using embedded coordinates as $\mathbf{Z}$ and a latent representation of, e.g., the simulation state as $\mathbf{Z}_{KV}$, cross-attention acts as a conditional neural field (Wang et al., 2024b) as each query vector retrieves information from the latent representation to produce a prediction for a given query coordinate. It is worth noting that the neural field decoding via cross-attention operates on a point-wise basis and hence, is independent of the number of query points.
@@ 3 Anchored-Branched Universal Physics Transformers
We introduce Anchored-Branched Universal Physics Transformers (AB-UPT), which leverages two key modeling components: (i) the multi-branch architecture outlined in LABEL:{sec:method-multibranch} separates geometry encoding, surface-level simulation, and volume-level simulation into distinct branches that interact with each other; (ii) anchor attentiondescribed in Section˜3.2 enables modeling of complex physical dynamics by leveraging expressive dot-product attention (Vaswani et al., 2017) within a heavily reduced set of anchor tokens. To generate predictions for the full simulation mesh during inference, anchor attention employs cross-attention between (millions) of query tokens and the small amount of anchor tokens, which has linear inference complexity. Additionally, anchor attention can be seen as a conditional neural field, which allows creating arbitrary output resolutions, and enables the application of differential operators to enforce hard physical constraints such as divergence-free vorticity-field predictions (see Section˜3.3).
Figure 4: The Anchored-Branched Universal Physics Transformers model is a multi-branch Transformer neural operator using three distinct branches: a surface, geometry, and volume branch. The physics blocks within each branch share parameters and interact through cross-attention. To handle problems at industry-scale, we select a subset of $M$ points from the input point cloud. We call these points anchor tokens, and they serve as the keys and values for attention computation. This approach significantly reduces the complexity of attention, making it possible to scale our model to industry-standard simulation meshes. Query positions define where physical quantities are predicted. For simplicity, we assume anchors and queries are disjoint subsets of all positions, enabling evaluation at anchor locations. We will demonstrate in Section 4.6 that anchors can also be placed arbitrarily (e.g., on a regular grid), in which case they may fall in regions where evaluating physical quantities is not physically meaningful (e.g., points inside a car geometry).
@@ 3.1 Multi-branch architecture
Our model builds on the Universal Physics Transformer (UPT(Alkin et al., 2024a) and multi-branch Transformers for neural operators (Alkin et al., 2024b; Esser et al., 2024). Multi-branch Transformers consist of multiple branches (which can share parameters) that each handle different data streams or modalities. We define a geometry branch, a surface branch, and a volume branch (see Figures˜3 and 4 for an overview). Those branches interact differently in respective encoding, physics, and decoding blocks, and allow us to enable scalability to high-resolution outputs via the new anchor attention. We distinguish the geometry, surface, and volume branches by the superscripts g{{g}}, s{{s}}, and v{{v}}, respectively.
@@
$\mathcal{S}:\mathbf{X}^{{g}}\in\mathcal{M}\xrightarrow[]{\text{Embed}}\mathbf{H}^{{g}}\in\mathbb{R}^{N^{{g}}\times d}\xrightarrow[]{\text{Supernode pooling}}\mathbf{S}\in\mathbb{R}^{S\times d},$ (4)
@@
$\mathcal{G}:\mathbf{S}\xrightarrow[]{\text{Attention}}\mathbf{Z}^{{g}}\in\mathbb{R}^{S\times d}.$ (5)
@@
$\mathcal{E}^{{s}}:\mathbf{X}^{{s}}\xrightarrow[]{\text{Embed}}\mathbf{H}^{{s}}\xrightarrow[]{\text{MLP}}\mathbf{Z}^{{s}}\in\mathbb{R}^{N^{{s}}\times d},$ (6a)
$\mathcal{E}^{{v}}:\mathbf{X}^{{v}}\xrightarrow[]{\text{Embed}}\mathbf{H}^{{v}}\xrightarrow[]{\text{MLP}}\mathbf{Z}^{{v}}\in\mathbb{R}^{N^{{v}}\times d}.$ (6b)
@@
$\mathcal{P}^{{s}}$ $:\big(\mathbf{Q}(\mathbf{Z}^{{s}}),\mathbf{K}(\mathbf{Z}^{{g}}),\mathbf{V}(\mathbf{Z}^{{g}})\big)\xrightarrow[]{\text{Attention}}\mathbf{Z}_{0}^{{s}},$ (7a)
$\mathcal{P}^{{v}}$ $:\big(\mathbf{Q}(\mathbf{Z}^{{v}}),\mathbf{K}(\mathbf{Z}^{{g}}),\mathbf{V}(\mathbf{Z}^{{g}})\big)\xrightarrow[]{\text{Attention}}\mathbf{Z}_{0}^{{v}},$ (7b)
@@
The surface and volume branch consists of $K$ physics blocks. Each physics block $k$, where $k\in\{1,\dots,K\}$, consists of a self-attention and cross-attention block, which share parameters between the surface and volume branches. Each physics block starts with a self-attention block, where the queries, keys, and values are all derived from the output $\mathbf{Z}_{k-1}$ of the preceding physics block k1k-1, where $\mathbf{Z}_{0}$ from Equation˜7 is the input to the first block. Next is a cross-attention block, enabling surface-volume interaction. For the surface branch, the queries originate from the previous block within the surface branch. However, the keys and values come from the volume branch $k$. For the volume branch, however, it is reversed: the queries come from the volume branch, and the keys and values from the surface branch. These cross-attention layers allow for intricate interactions between the surface and volumetric fields. Formally, physics block $k$ for the surface and volume branch, $\mathcal{B}^{{s}}_{k}$ and $\mathcal{B}^{{v}}_{k}$, are defined by:
$\mathcal{B}^{{s}}_{k}$ $:\big\{\mathbf{Q}(\mathbf{Z}^{{s}}_{k-1}),\mathbf{K}(\mathbf{Z}^{{s}}_{k-1}),\mathbf{V}(\mathbf{Z}^{{s}}_{k-1})\big\}\xrightarrow[]{\text{Attention}}\mathbf{Z}^{\prime{s}}_{k}\xrightarrow[]{}\big\{\mathbf{Q}(\mathbf{Z}^{\prime{s}}_{k}),\mathbf{K}(\mathbf{Z}^{\prime{v}}_{k}),\mathbf{V}(\mathbf{Z}^{\prime{v}}_{k})\big\}\xrightarrow[]{\text{Attention}}\mathbf{Z}^{{s}}_{k+1},$ (8a)
$\mathcal{B}^{{v}}_{k}$ $:\big\{\mathbf{Q}(\mathbf{Z}^{{v}}_{k-1}),\mathbf{K}(\mathbf{Z}^{{v}}_{k-1}),\mathbf{V}(\mathbf{Z}^{{v}}_{k-1})\big\}\xrightarrow[]{\text{Attention}}\mathbf{Z}^{\prime{v}}_{k}\xrightarrow[]{}\big\{\mathbf{Q}(\mathbf{Z}^{\prime{v}}_{k}),\mathbf{K}(\mathbf{Z}^{\prime{s}}_{k}),\mathbf{V}(\mathbf{Z}^{\prime{s}}_{k})\big\}\xrightarrow[]{\text{Attention}}\mathbf{Z}^{{v}}_{k+1}.$ (8b)
@@ 3.2 Anchor attention
High-fidelity numerical simulations in industry often rely on extremely fine mesh discretizations, sometimes comprising hundreds of millions of cells (Ashton et al., 2024b). In practice, representing such large point clouds (meshes) as inputs to surrogate models becomes computationally challenging. In particular, Transformer models (Vaswani et al., 2017)—the workhorse of many modern deep learning architectures—capture dependencies among all $N$ input tokens via self‐attention:
@@
To mitigate this problem, we introduce the concept of anchor attention, which splits the discretized input mesh (i. e., the point cloud) into two parts: (i) a small subset $\mathbf{A}$ of $M$ anchor tokens is uniformly sampled from the $N$ input tokens and used to derive the key matrix $\mathbf{K}(\mathbf{A})$ and value matrix $\mathbf{V}(\mathbf{A})$ for the attention mechanism; (ii) the query matrix $\mathbf{Q}(\mathbf{Z})$, which is derived from all $N$ input tokens, then cross-attend to the anchor key and value matrices. Intuitively, this formulation splits the $N$ input tokens into $M$ anchor tokens (used as queries, keys, and values in the attention) and NMN-M query tokens (used only as queries in the attention). It is equivalent to self-attention within the $M$ anchor tokens and cross-attention between NMN-M query tokens and the $M$ anchor tokens. Formally, anchor attention is defined as follows:
@@
For notational simplicity, we assume that anchor and query tokens come from the same set of points (e.g., positions from a CFD mesh). However, anchor and query tokens can also be obtained from completely different sets. We will demonstrate this in Section 4.6 where anchor tokens are sampled from a CAD geometry (surface anchors) or a regular grid (volume anchors) and the query tokens are obtained from the CFD mesh (both surface and volume query tokens).
@@
This design allows us to focus on problem scale (c. f. Figure˜2): by choosing the number of anchor tokens $M$, we directly control the model’s computational capacity, while enabling pointwise predictions at arbitrary points in the continuous output domain. These anchor tokens span both the surface and volume mesh, and their selection is crucial for capturing problem-specific structure. Increasing $M$ enhances contextual information available at each query location without changing the amount of model parameters, thereby improving performance up to a saturation point (see Figure˜5).
@@
Importantly, the NMN-M query tokens only need to compute attention with a much smaller set of $M$ key-value pairs (derived from the anchor tokens) compared to full self-attention, which computes all $N^{2}$ pairs. Anchor attention therefore reduces the computational complexity of the attention layers from $\mathcal{O}\left(N^{2}\right)$ to $\mathcal{O}\left(M^{2}+(N-M)M\right)=\mathcal{O}\left(MN\right)$, where $M^{2}$ is the self-attention between the $M$ anchor tokens and (NM)M(N-M)M is the cross-attention from NMN-M query tokens to $M$ anchor tokens. This is crucial because surrogate models often need to be evaluated at a large number of input points (e. g., to obtain accurate drag/lift coefficient predictions as shown in Figure 1).
@@
Additionally, since there are no interactions between query tokens, an arbitrary number of query tokens can be used per forward pass. Therefore, query tokens can be processed in chunks, enabling constant memory consumption by using a constant number of queries. When processing all $N$ input tokens at once (end-to-end inference), the memory complexity of anchor attention is $\mathcal{O}(N)$ (we use an efficient attention implementation (Dao et al., 2022)). However, as $N$ can be >> 100 million, it would still run into out-of-memory errors on a single GPU. By chunking the NMN-M query tokens into chunks of size $C$, memory complexity reduces to $\mathcal{O}(M+C)$, i.e., constant memory consumption. We refer to this methodology as chunked inference (c. f. inference memory in Figure˜1). Moreover, in chunked inference, the key-value pairs of anchor tokens can be cached when calculating them in the first chunked forward pass to omit recalculating them for subsequent chunks (anchor tokens remain constant for all chunks). However, as $M\ll N$, this is often negligible.
@@
Finally, by treating the anchors as conditioning set for the queries, anchor attention turns the architecture into a conditional field model: the field values at any query location are predicted from the query location itself and the context provided by the anchors, but independently from any other query location. This has two important benefits. First, this enables creating predictions at arbitrary query positions, which allows AB-UPT to predict at arbitrary resolution. Many simulation insights can be obtained from much lower resolutions than necessary for numeric simulation (e.g., drag/lift coefficients in Figure 6). Second, it allows us to impose physics‐based constraints directly into the model predictions, such as a divergence‐free condition for vorticity (see Section˜3.3). In the context of Neural Processes, the concept of conditioning the output prediction on a (small) subset of context has been explored in (Feng et al., 2023; Ashman et al., 2024).